OpenStudy (anonymous):

@jigglypuff314 Which of the following is not a property of a trapezoid?

OpenStudy (anonymous):

A.One pair of Parallel Lines B.Diagonals Are Congruent C.Length of the mid-segment is one-half the sum of the bases. D.Adjacent Angles are supplementary.

OpenStudy (anonymous):

Its either C or D.

jigglypuff314 (jigglypuff314):

D would be correct :)

OpenStudy (anonymous):

A science teacher has a supply of 50% sugar solution and a supply of 80% sugar solution. How much of each solution should the teacher mix together to get 150 mL of 60% sugar solution? a. 70 mL of the 50% solution and 35mL of the 80% solution. n. 35mL of the 50% solution and 70mL of the 80% solution. c. 70mL of the 50% solution and 70mL of the 80% solution. d. 35mL of the 50% solution and 35mL of the 80% solution.

jigglypuff314 (jigglypuff314):

sorry @firsst_queenx4 I'm not good at that sort of stuff :/

OpenStudy (anonymous):

wtf is this

OpenStudy (anonymous):

this is mah post

OpenStudy (anonymous):

mhm she backed off xD

OpenStudy (anonymous):

@Hello_Im_Nick , umm? that is Math dumbass & backed off what? you didn't say anything insulting, sooo....? You see numbers don't you? & thanks anyways @jigglypuff314

jigglypuff314 (jigglypuff314):

5x - 32 = 2x + 61

OpenStudy (anonymous):

xD Dumbass? How am I a dumbass? All I was saying was to get off my post....

OpenStudy (anonymous):

x=31 :D

jigglypuff314 (jigglypuff314):

10x + 7 = 2x + 41

OpenStudy (anonymous):

x=4.25 :D

OpenStudy (anonymous):

Would it be 42 Degrees?

jigglypuff314 (jigglypuff314):

yep :)

jigglypuff314 (jigglypuff314):

and for the previous, x = 5

OpenStudy (anonymous):

No.? x=4.25

jigglypuff314 (jigglypuff314):

oh wait, yeah, sorry >,<

jigglypuff314 (jigglypuff314):

5x + 13 + 3x - 5 = 180

OpenStudy (anonymous):

21.5?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):